# Rui Eduardo Lopes, Student Member, IEEE, Susana Sargento, Senior Member, IEEE

### v1

@@ III-D Inter- and intra-vehicular communication
Vehicular systems need to communicate both with other systems within and outside the vehicle. While most intra-vehicular communication is expected to be through wired physical layers, this does not have to be the norm. It is impractical to consider wired physical layer for inter-vehicular contexts.
Within inter-vehicular contexts, the Vehicle to Everything (V2X) paradigm, is of increasing importance considering the advent of both Vehicle to Vehicle (V2V) scenarios and smart cities.
The usage of cloud infrastructure is also an integrated part of the SDV paradigm, requiring Vehicle to Infrastructure (V2I) wireless communications. For example, CARIAD, the initiative from the Volkswagen group towards SDVs, launched an uniform application store for all brands within the Volkswagen group [27], that indicates a strong cloud presence within its ecosystem.
While intra-vehicular systems tend to be interconnected with CAN, LIN, FlexRay, or Media Oriented Systems Transport (MOST), they are unsuitable to the higher volume of data and interconnections of centralized architectures. Therefore, existing solutions, such as Ethernet, if associated with determinism mechanisms (further discussed in section III-E), can provide the required time-critical high throughput to both Input/Output (I/O) and control data streams [14].
@@
TABLE III: Overall comparison of main intra vehicular wired buses [14].
@@
[t] Characteristic / Bus LIN CAN FlexRay Ethernet Media 1 wire UTP 2/4 wires UTP Topology Bus Bus Star, bus Star Access control Master-slave Priority-based messages TDMA Switched point-to-point link Max. Baud Rate 19.2 Kb/s 1 Mb/s, up 16 Mb/sa(XL) 10 Mb/s 10 Mb/s, 100 Mb/s, 1000+ Mb/s MTU (Bytes) 8 8/64 254 1472 (UDP), 1460 (TCP), 9000 (Jumbo) Max nodes 16 (1 master) 16 (500 Kb/s) 22 (bus), 22/64 (star), 64 (mixed) 256 to theoretically limited Security Modules None SecOC SecOC MACsec, IPsec, TLS/DTLS, SecOC
@@
CAN has a limitation of 1 megabit per second of date rate, while CAN Flexible Data Rate (CAN-FD) raises the practical maximum date to 6 megabits per second with a higher 64 byte data field, while keeping the limitations on number of nodes and cable length [14].
While legacy buses like CAN or LIN can and will continue to be on edge intra vehicular systems, such as smaller more specific ECUs, sensors and actuators, the long term objective is the edges nodes to be also connected through Ethernet [65]. In the meantime, automakers aim at integrating legacy CAN or LIN based components by integrating in zonal ECUs CAN controllers or gateways [66].
While these limitations could be overcome in the past with a static analysis of the overall vehicular systems and ECUs, the growing number of features, specially data-intensive features related to (semi) autonomous driving, required an increase in the volume of the overall wiring warness both in terms of number of wires and of complexity (with heavy at least 40-50 kg [^43] Source: https://semiengineering.com/shedding-pounds-in-automotive-electronics/.) now commonplace).
@@
In addition, the CAN — and its variations, such as LIN [^44] Lower cost bus with just 2 wires (including ground) at the cost of lower throughput. — is based on broadcasting messages. This contributes towards the biggest flaw of the CAN — its fragility to the injection of counterfeit messages [^45] Example at https://kentindell.github.io/2023/04/03/can-injection/.. While filtering of messages (at the receiver level) is a possibility, it requires additional implementation on software (increasing the overall development complexity), and does not guarantee the CAN bus capability of handling and delivery other, genuine, messages is not affected [4].
Ethernet, within the context of vehicular networks, is being more and more considered as the backbone connection between high performance ECUs, with gateways to legacy buses like CAN. Several physical layer specifications can be considered, including IEEE 100BASE-T1, 1000BASE-T1, and 1000BASE-RH [67, 68]. Possible Ethernet backbone upgrades to ensure future scalability are also being explored, including 10BASE-T1S or multi-Gb links.
As depicted in fig. 8, there are multiple iterations of Ethernet standards, i.e. there’s not a single cable and solution for all scenarios. Automakers shall take into consideration the characteristics of the cables (such as width, weight or resistance to heat and vibrations)— and not only the theoretical bandwidth — in order to choose the best ones to interconnect vehicular ECUs.
@@
To this goal, major automakers and their providers have created the OPEN Alliance (One-Pair Ether-Net) [^46] Available at https://www.opensig.org/. initiative. Canonical also defends E/E architectures which combines and consolidates ECUs and interconnects them using an Ethernet network [6]. Such solutions could also benefit from industry grade protocols, such as PROFINET [^47] Details at https://www.profibus.com/technology/profinet/..
@@
Other alternatives Unlike CAN, FlexRay manages the medium with Time-Division Multiple Access (TDMA). FlexRay has a communication cycle with an explicit stage (the Network Idle Time) necessary to maintain synchronization between node clocks.
FlexRay has the unique ability to sync up nodes on a network without an external synchronization clock signal. To do so, it uses two special types of frames: Startup Frames, and Sync Frames. Like CAN, if not properly configured, may create starvation for some slots of data [72]. This protocol considers a eXtended Markup Language (XML) file format, the FIeld Bus EXchange, which contains configuration details on a network, including signal definitions, bit-level encoding of signals, information on network topology (e.g. information on the ECUs) and configurations (such as baud rates and timings).
MOST is an alternative. The MOST specification defines “the protocol, hardware and software layers necessary to allow for the efficient and low-cost transport of control, real-time and packet data” using “different physical layers (fiber optic, UTP, coax)” and supporting 25, 50 and 150 Mbps data transfer speeds [73].
Vehicular nodes can use different wireless access technologies to communicate with other road elements, including vehicles (V2V), pedestrians ( Vehicle to Pedestrian (V2P)), and infrastructure (V2I), creating a Vehicular Ad-hoc Network (VANET), a sub-type of a Mobile Ad-hoc Networks. These technologies can be classified based on their communication range [74], with a summary of the main characteristics of the most used access technologies described in table IV.
TABLE IV: Example of wireless vehicular communications access technologies and main characteristics.
@@
Vehicular targeted standards exist. Cellular Vehicle-to-Everything (C-V2X) is cellular based and aims at integrating cellular technologies within V2V scenarios. Based on cellular standards (5.9 GHz band), it also allows direct device-to-device (ad-hoc) communication with low latency and high reliability, but in short ranges (±1 km), with low cost/free of cost. C-V2X also support device-to-network, enabling cloud services to be part and parcel of the end-to-end solution.
ITS-G5 is based on IEEE 802.11p for physical and MAC layer. It requires existence of road infrastructure — Road Side Units — and that vehicles are equipped with On-Board Units. It suffers from problems on handling handovers and scaling to many nodes simultaneously. It is also vulnerable to packet collisions due to concurrent transmissions, and packet collisions due to hidden vehicular nodes.
Existing literature shows that C-V2X offers performance advantages over ITS-G5 in terms of its additional link budget, higher resilience to interference, and better non-line-of-sight capabilities [83].
Radio Access Technology (RAT) evolutions, such as the IEEE 802.11bd and New Radio for C-V2X can supplement today’s vehicular networks in supporting higher bandwidth and QoS wireless communications. IEEE 802.11bd are more suitable for supporting more V2X applications efficiently, with better throughput and low transmission latency. IEEE 802.11bd adapts the physical and MAC layer techniques introduced in IEEE 802.11n/ac/ax, while ensuring interoperability and backwards compatibility with 802.11p. This implies that 802.11bd and 802.11p devices must be able to communicate with each other while operating on the same channel [84]. IEEE 802.11bd also considers re-transmissions, Dual Carrier Modulation and operates also in the mmWave spectrum.
@@
III-E Deterministic inter- and intra-vehicular communication
The usage of deterministic communications and networks that support such communications is vital when considering vehicular services that are somehow critical. Determinism “is the quality or state of being determined” [^48] Definition per Merriam-Webster dictionary, available at https://www.merriam-webster.com/dictionary/determinism., i.e. the quality of being firmly resolved. Within networks this concept can be regarded as the delivery of units of information — either abstracted in frames or packets or other abstraction — in a fully resolved manner in several domains, such as temporal or reliability, in all the steps of the frame/packet from the producer to the consumer(s) of information.
Regarding the temporal determinism, within the exchange of a unit of data between two network entities one may consider the main steps of a packet switching as presented in fig. 9.
@@ III-E Deterministic inter- and intra-vehicular communication
Each step introduces delay that, if the network is not properly configured, can be unbounded and potentially even random. Even if high bandwidth is a possibility, transmission, and queuing delays still exist, affecting end-to-end delay [83, 85].
High performance and high precision of services running through a communication network should consider not only techniques to decrease the overall end-to-end latency of data from/to services — for example by leveragging the edge computing concept [88, 89] — but also techniques to ensure the latency requirements and real-time hard constraints are followed. Sets of standards such as TSN [^49] Details at https://1.ieee802.org/tsn/. and Deterministic Networks [^50] Details at https://datatracker.ietf.org/wg/detnet/about/. can ensure that any critical messages from/to services arrive with high reliability and that arrive at the exact supposed moment [90].
Current solutions to tackle these issues within the TSN and Deterministic Networks are further presented in section III-E. The next subsections present the state-of-the-art on the main initiatives and sets of standards towards determinism in networks in lower Open Systems Interconnection (OSI) layers. These initiatives started for wired mediums such as Ethernet but can also be extended to wireless contexts.
III-E1 In wired context
Methods towards achieving deterministic communications within a wired context — such as the connectivity between vehicular ECUs — are being standardised, with two major initiatives to be considered: the Institute of Electrical and Electronics Engineers (IEEETSN set of standards, operating at OSI layer 2; and Internet Engineering Task Force (IETF) Deterministic Networking (DetNet) operating at OSI layer 3. The next subsections will focus on both initiatives that shall be seen not as competitors but as complementary initiatives.
Deterministic Networking (IETF DetNet)  The IETF Deterministic Networking set of standards aim at defining deterministic — i.e. with bounded latency, loss and jitter — paths for the packets [91]. Those standards operate at layer 3 on a SDN approach, where one or more controllers are responsible for defining the nodes behavior, to warrant [92]:
@@ III-E1 In wired context
Table V provides an overview on the Request for Comments (RFC) standards and documents that define the goals, use cases and building blocks behavior.
TABLE V: Main DetNet standards, specifications and documents.
@@
RFC 9016 YANG model (draft [93]) Flow information model Modelling of the configuration information for configuration of DetNet building blocks.
@@
The DetNet building blocks operate at the Internet Protocol (IP) layer and use lower-layer technologies such as MPLS and TSN (further discussed in section III-E1). These building blocks can and should be controlled based on observable metrics (hops, their state, and end-to-end latency), and on fault detection/prediction and identification [92].
Two open questions remain: (1) whether DetNet protocols will include a way for an application to communicate expectations regarding such timing properties to the network and (2) if so, whether those properties would likely have a material effect on network performance as a result [94].
Time Sensitive Networking (IEEE TSN) The Time Sensitive Network group presents an effort to solve transmission, and queuing delays issues, by adding to Ethernet time characteristics, performance, reliability, and security. Mechanisms include time synchronization (gPTP [86]), replication and elimination for reliability (e.g. Frame Replication and Elimination for Reliability (FRER)) [95], and bounded low latency (such as credit-based [96] or time aware shaping algorithms [97]) that can provide the necessary real-time performance and bounding of the latencies [98, 99, 100].
@@
TABLE VI: Main TSN standards [103] (partially adapted).
@@
In conjunction to DetNet, but operating at OSI layer 2 — and originally on top of wired Ethernet networks — the IEEE 802.1 Time Sensitive Network set of standards provide mechanisms for [103]:
  • Time synchronization: applications and other network mechanisms require a precise (down to nanosecond accuracy) synchronization between the nodes of the network. Protocols such as Precision Time Protocol (PTP[86] can provide such precision within wired and wireless networks. Synchronization should have into mind the physical layer latencies and their variations since they can affect the propagation delay measurement and the overall sync time accuracy. IEEE 802.3cx proposes a mechanism to report the physical layer delays in a dynamic way [104].
  • Replication and elimination for reliability: improve the reliability by ensuring frames of information are replicated and sent through disjoint paths. This replication needs additional mechanisms to ensure repeated frames are ignored. Protocols such as FRER [95] can manage this replication.
  • Resource management: Mechanisms to ensure the configuration of TSN talker, listener and of equipment between both.
  • Bounded latency or bandwidth: ensure the bandwidth or latency and its jitter is deterministic by ensuring transmission happens at a specific time. Traffic shaping can provide the necessary real-time performance and bounding of the latencies or bandwidth [98, 99, 100], with shapers such as:
    • Credit Based Shaper (CBS[96] based on regulating the transmission rate of frames to smooth out traffic, reduce bursts and ensure bandwidth for certain traffic classes, by delaying the transmission of some frames. When the queue contains a frane to be transmitted, a credit counter is checked and the frame is only transmitted if the counter is non-negative (and if it is, is held back until is non-negative). This counter accumulates credits when the transmission queue is idle, and consumes credits when frames are transmitted, with the rate of accumulation and consumption configurable via parameters (idle slope and send slope) [^51] More details at https://inet.omnetpp.org/docs/showcases/tsn/trafficshaping/
      creditbasedshaper/doc/index.html
    • Time Aware Shaper (TAS[97] that works by associating the data to traffic classes and the traffic classes to gates that open and close at specific times and for specific intervals of time, ensuring no unwanted competition between traffic classes that should not compete for transmission, therefore eliminating the jitter associated with that competition [105].
    • Asynchronous Traffic Shaping (ATS), which is based on the leaky bucket principle, with transmission allowed when credit is greater than the size of the frame [102].
@@
These three models are based on the existence of two entities — the CNC and the CUC. The CNC defines the overall network schedule and is the main controller of the network within the network bridges, dispatching configuration with protocols such as Network Configuration Protocol (NETCONF). The CUC mediates the requests between end devices and their applications [107], collects all user requirements and interacts with the CNC to decide the topology and scheduling of the network. Figure 10 provides an overview of these interactions.
@@
TSN is not however the only standard at layer 2 for deterministic communications. Time-Triggered Ethernet (TTEthernet), a standard for industrial and avionics applications improved from the Avionics Full Duplex Switched Ethernet, can be a standard to consider [14]. While the concepts of deterministic timing of the transmission of frames is inherently the same, both differ in a number of aspects [108, 109]. TSN is a set of standards while TTEthernet is a specific Society of Automotive Engineers (SAE) AS6802 standard. In the former, the schedule is on entire traffic classes, while in latter is on individual frames. In addition, TSN supports lower-priority traffic using non-used bandwidth, while TTEthernet blocks other traffic even when there are no time-scheduled frames to transmit results in efficient use of bandwidth.
III-E2 In wireless context
While wired network access mediums have well defined / bounded error rates, based essentially on the electromagnetic noise ratio, wireless mediums are prone to the creation of unpredictable non-stochastic jitter and packet losses through additional issues and challenges, including [103]:

### v2

@@ 3.5 Intra-vehicular communications
Vehicular systems need to communicate with other systems within and outside the vehicle. While most intra-vehicular communication is expected to be through wired physical layers, this does not have to be the norm. Table 3 provides an overview of the main buses that may be used for intra-vehicular communications.
@@
Table 3: Overall comparison of main intra vehicular wired buses [19].
@@
[t] Characteristic / Bus LIN CAN FlexRay Ethernet Media 1 wire Unshielded Twisted Pair cable (UTP) 2/4 wires UTP Topology Bus Bus Star, bus Star, bus Access control Master-slave Priority-based messages TDMA Switched point-to-point link Max. Baud Rate 19.2 Kb/s 1 Mb/s, up 16 Mb/s (XL)a 10 Mb/s 10 Mb/s, 100 Mb/s, 1000+ Mb/s MTU (Bytes) 8 8/64 254 1472 (UDP), 1460 (TCP), 9000 (Jumbo) Max nodes 16 (1 master) 16 (500 Kb/s) 22 (bus), 22/64 (star), 64 (mixed) 256 to theoretically limited Security Modules None SecOC SecOC MACsec, IPsec, TLS/DTLS, SecOC
@@
In the past, these limitations could be addressed through static analysis of the overall vehicular systems and ECUs. However, the increasing number of features —particularly data-intensive functionalities related to semi-autonomous driving— has necessitated a significant expansion in the volume and complexity of the wiring harness. This increase involves not only a greater number of wires but also added complexity, leading to a substantial weight of at least 40-50 kg [^45] Source: semiengineering.com..
While legacy buses like CAN or LIN can and will continue to be on edge intra vehicular systems, such as smaller more specific ECUs, sensors and actuators, the long term objective is the edges nodes to be also connected through Ethernet [61]. Within the context of vehicular networks, Ethernet is being more and more considered as the backbone connection between high performance ECUs, with gateways to legacy buses like CAN. As depicted in fig. 10, there are multiple iterations of Ethernet standards, i.e. there’s not a single cable and solution for all scenarios. Automakers shall take into consideration the characteristics of the cables (such as width, weight or resistance to heat and vibrations)— and not only the theoretical bandwidth — in order to choose the best ones to interconnect vehicular ECUs. Several physical layer specifications can be considered, including IEEE 100BASE-T1, 1000BASE-T1, and 1000BASE-RH [62, 63]. Possible Ethernet backbone upgrades to ensure future scalability are also being explored, including 10BASE-T1S or multi-Gb links.
@@
Towards the goal of support Ethernet on the vehicles, major automakers and their providers have created the OPEN Alliance (One-Pair Ether-Net) [^46] At opensig.org. initiative. Canonical also defends E/E architectures which combines and consolidates ECUs and interconnects them using an Ethernet network [3]. Such initiatives could also benefit from the resilience proven by industry grade protocols, such as PROFINET [^47] At profibus.com..
@@
Alternatives to Ethernet include FlexRay and Media Oriented Systems Transport (MOST). Unlike CAN, FlexRay has the unique ability to sync up nodes on a network without an external synchronization clock signal by using two special types of frames [68], being a possible alternative to Ethernet with Deterministic mechanisms (more on section 3.7). The MOST specification defines “the protocol, hardware and software layers necessary to allow for the efficient and low-cost transport of control, real-time and packet data” using “different physical layers (fiber optic, UTP, coax)” and supporting 25, 50 and 150 Mbps data transfer speeds [69], still under the theoretical maximus of Ethernet.
3.6 Inter-vehicular communications
For inter vehicular communications, it is impractical to consider wired physical layer. Within these context, the Vehicle to Everything (V2X) paradigm is of increasing importance considering the advent of both Vehicle to Vehicle (V2V) scenarios and smart cities. The usage of cloud infrastructure is also an integrated part of the SDV paradigm, requiring Vehicle to Infrastructure (V2I) wireless communications. For example, CARIAD, the initiative from the Volkswagen group towards SDVs, launched an uniform application store for all brands within the Volkswagen group [25], that indicates a strong cloud presence within its ecosystem.
Vehicular nodes can use different wireless access technologies to communicate with other road elements, including vehicles, V2V; pedestrians, Vehicle to Pedestrian (V2P); and infrastructure, V2I; creating a Vehicular Ad-hoc Network (VANET), a sub-type of a Mobile Ad-hoc Networks. These technologies can be classified based on their communication range [70], with a summary of the main characteristics of the most used access technologies described in table 4.
Table 4: Example of wireless vehicular communications access technologies and main characteristics.
@@ 3.6 Inter-vehicular communications
Vehicular targeted standards include Cellular Vehicle-to-Everything (C-V2X) and ITS-G5. C-V2X is cellular based and aims at integrating cellular technologies within V2V scenarios. Based on cellular standards it also allows direct device-to-device communication with low latency and high reliability, but in short ranges (±1 km), with low cost associated with the possible necessary hardware changes. C-V2X also support device-to-network, enabling cloud services to be part and parcel of the end-to-end solution. On the other hand, ITS-G5 is based on IEEE 802.11p for physical and MAC layer. It requires existence of road infrastructure — Road Side Units — and that vehicles are equipped with On-Board Units. Problems of this alternative includes handling handovers and scaling to many nodes simultaneously and vulnerability to packet collisions due to concurrent transmissions, and packet collisions due to hidden vehicular nodes. Existing literature shows that C-V2X offers performance advantages over ITS-G5 in terms of its additional link budget, higher resilience to interference, and better non-line-of-sight capabilities [79].
Radio Access Technology (RAT) evolutions, such as the IEEE 802.11bd and New Radio for C-V2X can supplement today’s vehicular networks in supporting higher bandwidth and QoS wireless communications. IEEE 802.11bd are more suitable for supporting more V2X applications efficiently, with better throughput and low transmission latency, by adapting the physical and MAC layer techniques introduced in IEEE 802.11n/ac/ax, while ensuring interoperability and backwards compatibility with 802.11p. [80].
3.7 Deterministic intra and inter-vehicular communications
@@
Services running through a communication network and with requirements of high performance and high precision should consider not only techniques to decrease the overall end-to-end latency of data from/to services — for example by leveragging the edge computing concept [83, 84] — but also techniques to ensure the latency requirements and real-time hard constraints. The usage of deterministic communications and networks that support such communications are vital when considering vehicular services that are somehow critical. Determinism, per the Merriam-Webster dictionary, “is the quality or state of being determined”, i.e. the quality of being firmly resolved. Within networks this concept can be regarded as the delivery of units of information — either abstracted in frames or packets or other abstraction — in a fully resolved manner in several domains, such as temporal or reliability, in all the steps of the frame/packet from the producer to the consumer(s) of information.
Regarding the temporal determinism, within the exchange of a unit of data between two network entities one may consider the main steps of a packet switching as presented in fig. 11. Each step introduces delay that, if the network is not properly configured, can be unbounded and potentially even random. Even if high bandwidth is a possibility, transmission, and queuing delays still exist, affecting end-to-end delay [79, 81].
@@
Sets of standards such as TSN [^48] At 1.ieee802.org/tsn. and Deterministic Networks [^49] At datatracker.ietf.org/wg/detnet. can ensure that high reliability and time precision for each step of the packet switching process [86]. Current solutions to tackle these issues within the TSN and Deterministic Networks are further presented in this section. The next subsections present the state-of-the-art on the main initiatives and sets of standards towards determinism in networks in lower Open Systems Interconnection (OSI) layers. These initiatives started for wired mediums such as Ethernet but can also be extended to wireless contexts.
3.7.1 In wired context
Methods towards achieving deterministic communications within a wired intra-vehicular context (such as the connectivity between vehicular ECUs) are being standardised, with two major initiatives to be considered: a) the Institute of Electrical and Electronics Engineers (IEEETSN set of standards, operating at OSI layer 2; and b) the Internet Engineering Task Force (IETF) Deterministic Networking (DetNet) operating at OSI layer 3. Both initiatives shall be seen not as competitors but as complementary initiatives.
The IETF Deterministic Networking set of standards aims at defining deterministic — i.e. with bounded latency, loss and jitter — paths for the packets [87]. Those standards operate at layer 3 and can be supported by a Software-Defined Network (SDN) approach, where one or more controllers are responsible for defining the nodes behavior, to warrant [88]:
  • 1.
@@ 3.7.1 In wired context
Table 5 provides an overview on the Request for Comments (RFC) standards and documents that define the goals, use cases and building blocks behavior. The DetNet building blocks operate at the Internet Protocol (IP) layer and use lower-layer technologies such as MPLS and TSN. These building blocks can and should be controlled based on observable metrics (hops, their state, and end-to-end latency), and on fault detection/prediction and identification [88].
Table 5: Main DetNet standards, specifications and documents.
@@
RFC 9016 YANG model (draft) [89] Flow information model Modelling of the configuration information for configuration of DetNet building blocks.
@@
At a lower OSI layer, the Time Sensitive Network group presents an effort to solve transmission, and queuing delays issues. Table 6 provides an overview on the IEEE standards and documents that define the goals, use cases and building blocks behavior. TSN standards adds to Ethernet mechanisms to ensure time characteristics, performance, reliability, and security. Mechanisms include time synchronization (e.g. gPTP [82]), replication and elimination for reliability (e.g. Frame Replication and Elimination for Reliability (FRER)) [90], and bounded low latency (e.g. credit-based [91] or time aware shaping algorithms [92]) that can provide the necessary real-time performance and bounding of the latencies [93, 94, 95].
@@
In conjunction with DetNet, but operating at OSI layer 2 — and originally on top of wired Ethernet networks — the IEEE 802.1 TSN set of standards provide mechanisms for [98]:
  • 1.
    Time synchronization: applications and other network mechanisms require a precise (down to nanosecond accuracy) synchronization between the nodes of the network. Protocols such as Precision Time Protocol (PTP[82] can provide such precision within wired and wireless networks. Synchronization must take into account physical layer latencies and their variations, as these can impact propagation delay measurements and overall synchronization accuracy. IEEE 802.3cx proposes a mechanism to report the physical layer delays in a dynamic way [99].
  • 2.
    Replication and elimination for reliability: improve the reliability by ensuring that frames are replicated and sent through disjoint paths. This replication needs additional mechanisms to ensure repeated frames are ignored. Solutions such as FRER [90] can manage this replication.
  • 3.
    Resource management: Mechanisms to ensure the configuration of TSN talker, listener and of equipment between both and the necessary allocation of resources.
  • 4.
    Bounded latency or bandwidth: ensure the bandwidth or latency and its jitter is deterministic by ensuring transmission happens at a specific time.
Regarding bounded latency, traffic shaping can provide the necessary real-time performance and bounding of the latencies or bandwidth [93, 94, 95], with shapers such as:
  • 1.
    Credit Based Shaper (CBS): based on regulating the transmission rate of frames to smooth out traffic, reduce bursts and ensure bandwidth for certain traffic classes, by delaying the transmission of some frames. When the queue contains a frame to be transmitted, a credit counter is checked and the frame is only transmitted if the counter is non-negative (and if it is, is held back until is non-negative). This counter accumulates credits when the transmission queue is idle, and consumes credits when frames are transmitted, with the rate of accumulation and consumption configurable via parameters (idle slope and send slope) [^50] More details at inet.omnetpp.org. [91].
  • 2.
    Time Aware Shaper (TAS): works by associating the data to traffic classes and the traffic classes to gates that open and close at specific times and for specific intervals of time, ensuring no unwanted competition between traffic classes that should not compete for transmission, therefore eliminating the jitter associated with that competition [92, 100].
  • 3.
    Asynchronous Traffic Shaping (ATS): which is based on the leaky bucket principle, with transmission allowed when credit is greater than the size of the frame [97].
Table 6: Main TSN standards [98] (partially adapted).
@@
  • 1.
    A fully distributed model where all the end stations communicate their requirements — through protocols such as Stream Reservation Protocol (SRP) or Resource Allocation Protocol (RAP).
  • 2.
    A centralized network/distributed user model, where an entity, the CNC, gather the network capabilities.
  • 3.
    A fully centralized model, where the CNC is helped by the CUC to ensure proper network equipment configuration.
These three models are based on the existence of two entities — the Centralized Network Controller (CNC) and the Centralized User Configuration (CUC). The CNC defines the overall network schedule and is the main controller of the network within the network bridges, dispatching configuration with protocols such as Network Configuration Protocol (NETCONF). The CUC mediates the requests between end devices and their applications [102], collects all user requirements and interacts with the CNC to decide the topology and scheduling of the network. Figure 12 provides an overview of these interactions.
@@
TSN is not however the only standard at layer 2 for deterministic communications. Time-Triggered Ethernet (TTEthernet), a standard for industrial and avionics applications improved from the Avionics Full Duplex Switched Ethernet, is an alternative[19]. While the concepts of deterministic timing of the transmission of frames is inherently the same, both differ in a number of aspects [103, 104]. TSN is a set of standards while TTEthernet is a specific SAE AS6802 standard. In the former, the schedule is on entire traffic classes, while in latter is on individual frames. In addition, TSN supports lower-priority traffic using non-used bandwidth, while TTEthernet blocks other traffic even when there are no time-scheduled frames to transmit results in a less efficient use of bandwidth.
3.7.2 In wireless context
While wired network access mediums have well defined / bounded error rates, based essentially on the electromagnetic noise ratio, wireless mediums (such as inter vehicular communications are prone to the creation of unpredictable non-stochastic jitter and packet losses through additional issues and challenges, including [98]:
  • 1.